In a one-tailed hypothesis test (upper tail), the test statistic is determined to be 2.5. the p-value for this test is

a. .0228

b. .0062

c. .0056

d. .5228
Answer #1

Solution

Upper tailed z value is 2.5

P(z > 2.5) = 1 - P(z < 2.5) = 1 - 0.9938 = 0.0062

P-value = 0.0062